# Multi-agent cooperation through learning-aware policy gradients

### v1

@@ Abstract
Self-interested individuals often fail to cooperate, posing a fundamental challenge for multi-agent learning. How can we achieve cooperation among self-interested, independent learning agents? Promising recent work has shown that in certain tasks cooperation can be established between learning-aware agents who model the learning dynamics of each other. Here, we present the first unbiased, higher-derivative-free policy gradient algorithm for learning-aware reinforcement learning, which takes into account that other agents are themselves learning through trial and error based on multiple noisy trials. We then leverage efficient sequence models to condition behavior on long observation histories that contain traces of the learning dynamics of other agents. Training long-context policies with our algorithm leads to cooperative behavior and high returns on standard social dilemmas, including a challenging environment where temporally-extended action coordination is required. Finally, we derive from the iterated prisoner’s dilemma a novel explanation for how and when cooperation arises among self-interested learning-aware agents.
@@ 1 Introduction
From self-driving autonomous vehicles to personalized assistants, there is a rising interest in developing agents that can learn to interact with humans (Collins et al., 2024; Gweon et al., 2023), and with each other (Park et al., 2023; Vezhnevets et al., 2023). However, multi-agent learning comes with significant challenges that are not present in more conventional single-agent paradigms. This is perhaps best seen through the study of ‘social dilemmas’, general-sum games which model the tension between cooperation and competition in abstract form (von Neumann and Morgenstern, 1947). Without further assumptions, letting agents independently optimize their individual objectives on such games results in poor outcomes and a lack of cooperation (Tan, 1993; Claus and Boutilier, 1998).
@@
Optimal single-agent behavior can be formalized through the theory of stochastic optimal control on Markov decision processes (Bellman, 1957; Åström, 1965), which provides a basis for reinforcement learning (Sutton and Barto, 2018). The independent multi-agent learning problems that we study here are more complicated. First, for general-sum games, reaching an equilibrium point does not necessarily imply appropriate behavior because there can be many sub-optimal equilibria (Fudenberg and Levine, 1998; Shoham and Leyton-Brown, 2008). Second, the control problem an agent faces is non-stationary from its own viewpoint, because other agents themselves simultaneously learn and adapt (Hernandez-Leal et al., 2017). Centralized training algorithms sidestep non-stationarity issues by sharing agent information (Sunehag et al., 2017), but this transformation into a global learning problem is usually prohibitively costly, and impossible to implement when agents must be developed separately (Zhang et al., 2021).
@@
The above two fundamental issues have hindered progress in multi-agent reinforcement learning, and have limited our understanding of how self-interested agents may reach high returns when faced with social dilemmas. In this paper, we join a promising line of work on ‘learning awareness’ (Foerster et al., 2018a). The key idea behind such approaches is to take into account the learning dynamics of other agents explicitly, rendering it into a meta-learning problem (Schmidhuber, 1987; Bengio et al., 1990; Hochreiter et al., 2001). Foerster et al. (2018a) showed that two players could reach cooperation in the infamous iterated prisoner’s dilemma by using a simple form of anticipation of each other’s learning, whereas standard, independent learning would fail to produce cooperation.
@@
The present paper contains two main novel results on learning awareness in general-sum games. First, we introduce a new learning-aware reinforcement learning rule derived as a policy gradient estimator. Unlike existing methods (Foerster et al., 2018a, b; Xie et al., 2021; Balaguer et al., 2022; Lu et al., 2022; Willi et al., 2022; Cooijmans et al., 2023; Khan et al., 2024), it has a number of desirable properties: (i) it does not require computing higher-order derivatives, (ii) it is provably unbiased, (iii) it can model minibatched learning algorithms, and (iv) it is applicable to scalable architectures based on recurrent sequence policy models. This policy gradient rule significantly outperforms previous methods in the general-sum game setting. Notably, we show that it reaches cooperation for the first time in a challenging sequential social dilemma involving temporally extended actions (Leibo et al., 2017) that we adapted from the Melting Pot suite (Agapiou et al., 2023). Second, we analyze the iterated prisoner’s dilemma, a canonical model for studying cooperation among self-interested agents (Rapoport, 1974; Axelrod and Hamilton, 1981), and we discover a previously unknown mechanism for the emergence of cooperation through learning awareness. The mechanism we identify relies on heterogeneous groups of agents with various degrees of learning awareness, further reinforcing the importance of population heterogeneity in overcoming social dilemmas (Nowak and Sigmund, 1992; Santos et al., 2006).
@@ 2 Background and problem setup
We consider partially observable stochastic games (POSGs; Kuhn, 1953) consisting of a tuple $(\mathcal{I},\mathcal{S},\allowbreak\mathcal{A},P_{t},P_{r},P_{i},\mathcal{O},P_{o},\gamma,T)$ with =$\mathcal{I}=\{1,\ldots,n\}$ a finite set of $n$ agents, $\mathcal{S}$ the state space, $\mathcal{A}=\times_{i\in\mathcal{I}}\mathcal{A}^{i}$ the joint action space, $P_{t}(S_{t+1}\mid S_{t},A_{t})$ the state transition distribution, $P_{i}(S_{0})$ the initial state distribution, $P_{r}=\times_{i\in\mathcal{I}}P_{r}^{i}(R\mid S,A)$ the joint factorized reward distribution with $R=\{R^{i}\}_{i\in\mathcal{I}}$ and bounded rewards $R^{i}$, $\mathcal{O}=\times_{i\in\mathcal{I}}\mathcal{O}^{i}$ the joint observation space, $P_{o}(O_{t}\mid S_{t},A_{t-1})$ the observation distribution, $\gamma$ the discount factor, $t$ the time step, and $T$ the horizon. We use superscript $i$ to indicate agent-specific actions, observations and rewards, $-i$ to indicate all agent indices except $i$, and we omit the superscript for joint actions, observations and rewards. As agents only receive partial state information, they benefit from conditioning their policies $\pi^{i}(a^{i}_{t}\mid x^{i}_{t};\phi^{i})$ on the observation history $x^{i}_{t}=\{o^{i}_{k}\}_{k=1}^{t}$ (Åström, 1965; Kaelbling et al., 1998), with $\phi^{i}$ the policy parameters. Note that the observations can contain the agent’s actions on previous timesteps.
@@ 2.1 General-sum games and their challenges
We focus on general-sum games, where each agent has their own reward function, possibly different from those of other agents. Analyzing and solving general-sum games while letting every agent individually and independently maximize their rewards (a setting often referred to as “fully-decentralized reinforcement learning”; Albrecht et al., 2024) is a longstanding problem in the fields of machine learning and game theory. Alongside the difficulties of standard reinforcement learning, this setting comes with two additional significant issues that we briefly review below.
Non-stationarity of the environment.
In a general-sum game, each agent aims to maximize its expected return $J^{i}(\phi^{i},\phi^{-i})=\mathbb{E}_{P^{\phi^{i},\phi^{-i}}}\left[\sum_{t=1}^{T}\gamma^{t}R^{i}_{t}\right]$, with $P^{\phi^{i},\phi^{-i}}$ the distribution over environment trajectories $x_{T}$ induced by the environment dynamics, the policy $\pi^{i}(a^{i}\mid x^{i};\phi^{i})$ of agent $i$, and the policies $\pi^{-i}$ of all other agents. Importantly, the expected return $J^{i}(\phi^{i},\phi^{-i})$ does not only depend on the agent’s own policy, but also on the current policies of the other agents. As other agents are updating their policies through learning, the environment which includes the other agents is effectively non-stationary from a single agent’s perspective. Furthermore, the actions of an agent can influence this non-stationarity by changing the observation histories of other agents, on which they base their learning updates.
Equilibrium selection.
@@ 2.2 Co-player learning awareness
We aim to address the above two major challenges of multi-agent learning in this paper. Our work builds upon recent efforts that are based on adding a meta level to the multi-agent POSG, where the higher-order variable represents the learning algorithm used by each agent (Lu et al., 2022; Balaguer et al., 2022; Khan et al., 2024). In this meta-problem, the environment includes the learning dynamics of other agents. At the meta-level, one episode now extends across multiple episodes of actual game play, allowing the ‘ego agent’, $i$, to observe how its co-players, $-i$, learn, see Fig. 1. The goal of this meta-agent may be intuitively understood as that of shaping co-player learning to its own advantage.
@@
Following Lu et al. (2022), we first introduce a meta-game with a single meta-agent whose goal is to shape the learning of naive co-players to its advantage. This meta-game is defined formally as a single-agent partially observable Markov decision process (POMDP) $(\tilde{\mathcal{S}},\tilde{\mathcal{A}},\tilde{P}_{t},\tilde{P}_{r},\tilde{P}_{i},\tilde{\mathcal{O}},\tilde{\gamma},M)$. The meta-state consists of the policy parameters $\phi^{-i}$ of all co-players together with the agent’s own parameters $\phi^{i}$. The meta-environment dynamics represent the fixed learning rules of the co-players, and the meta-reward distribution represents the expected return $J^{i}(\phi^{i},\phi^{-i})$ collected by agent $i$ during an inner episode, with ‘inner’ referring to the actual game being played. The initialization distribution $\tilde{P}_{i}$ reflects the policy initializations of all players. Finally, we introduce a meta-policy $\mu(\phi^{i}_{m+1}\mid\phi^{i}_{m},\phi^{-i}_{m};\theta)$ parameterized by $\theta$, that decides the update to the parameter $\phi^{i}_{m+1}$ (the meta-action) to shape the co-player learning towards highly rewarding regions for agent $i$ over a horizon of $M$ meta steps. This leads to the co-player shaping problem
@@ 2.3 Single-level co-player shaping by leveraging sequence models
To proceed with this approach, we must first reformulate the meta-game. In particular, we must deal with a difficulty that is not present in single-agent meta reinforcement learning (e.g., Wang et al., 2016; Duan et al., 2017), which stems from the fact that co-players generally update their own policies based on multiple inner episodes (‘minibatches’), without which reinforcement learning cannot practically make progress. Here, we solve this by defining the environment dynamics over $B$ parallel trajectories, with $B$ the size of the minibatch of inner episode histories that co-players use to update their policies at each inner episode boundary (see Fig. 1A).
Batched co-player shaping POMDP.
@@
We define the batched co-player shaping POMDP $(\bar{\mathcal{S}},\bar{\mathcal{A}},\allowbreak\bar{P}_{t},\bar{P}_{r},\bar{P}_{i},\bar{\mathcal{O}},\bar{\gamma},M,B)$, with hidden states consisting of the hidden environment states of the $B$ ongoing inner episodes, combined with the current parameters $\phi^{-i}_{m}$ of all co-players; environment dynamics $\bar{P}_{t}$ simulating $B$ environments in parallel, combined with updating the co-player’s policy parameters $\phi^{-i}$, and resetting the environments at each inner episode boundary; initial state distribution $\bar{P}_{i}$ that initializes the co-player policies and initializes the environments for the first inner episode batch; and finally, an ego-agent policy $\bar{\pi}^{i}(\bar{a}^{i}_{l}\mid\bar{h}^{i}_{l};\phi^{i})$ parameterized by $\phi^{i}$, which determines a distribution over the batched action $\bar{a}^{i}_{l}=\{a^{i,b}_{l}\}_{b=1}^{B}$, based on the batched long history $\bar{h}^{i}_{l}=\{h^{i,b}_{l}\}_{b=1}^{B}$. We refer to each element of the latter as a long history $h^{i,b}_{l}$, with long time index $l$ running across multiple episodes, from $l=1$ until $l=MT$. It should be contrasted to the inner episode history $x_{t}^{i}$, which runs from $t=1$ to $t=T$ and thus only reflects the current (inner) game history.
@@
The POMDP introduced above suggests using a sequence policy $\bar{\pi}^{i}(\bar{a}^{i}_{l}\mid\bar{h}^{i}_{l};\phi^{i})$ that is aware of the full minibatch of long histories and which produces a joint distribution over all current actions in the minibatch. However, as we aim to use our agents not only to shape naive learners, but also to play against/with each other, we require a policy that can be used both in a batch setting with naive learners, and in a single-trajectory setting with other learning-aware agents. Within our single-level approach, we achieve this by factorizing the batch-aware policy $\bar{\pi}^{i}(\bar{a}^{i}_{t}\mid\bar{h}^{i}_{l};\phi^{i})$ into $B$ independent policies with shared parameters $\phi^{i}$,
$\bar{\pi}^{i}(\bar{a}^{i}_{l}\mid\bar{h}^{i}_{l};\phi^{i})=\prod_{b=1}^{B}\pi^{i}(a^{i,b}_{l}\mid h^{i,b}_{l};\phi^{i}).$ (2)
Thanks to the batched POMDP, we can now pose co-player shaping as a standard (single-level, single-agent) expected return maximization problem:
@@ 3 Co-agent learning-aware policy gradients
3.1 A policy-gradient for shaping naive learners
@@
$\nabla_{\phi^{i}}\bar{J}(\phi^{i})=\mathbb{E}_{\bar{P}^{\phi^{i}}}\left[\sum_{b=1}^{B}\sum_{l=1}^{MT}\nabla_{\phi^{i}}\log\pi^{i}(a^{i,b}_{l}\mid h^{i,b}_{l})\left(\frac{1}{B}\sum_{b^{\prime}=1}^{B}\sum_{k=l}^{MT}R_{k}^{i,b^{\prime}}\right)\right].$ (4)
We provide a proof in Appendix C. There are two important differences between COALA-PG and naively applying policy gradient methods to individual trajectories in a batch. (i) Each gradient term for an individual action $a_{l}^{i,b}$ takes into account the future return averaged over the whole minibatch, instead of the future return along trajectory $b$ (see Fig. 2). This allows taking into account the influence of this action on the parameter update of the naive learner, which influences all trajectories in the minibatch. (ii) Instead of averaging the policy gradients for each trajectory in the batch, COALA-PG accumulates (sums) them. This is important, as otherwise the learning signal would vanish in the limit of large minibatches. Intuitively, when a naive learner uses a large minibatch for its updates, the effect of a single action on the naive learner’s update is small, creating the need to sum the small effects together to obtain a policy gradient with a magnitude that does not depend on batch size. We will later show experimentally in Section 6 that correct treatment of minibatches critically affects reinforcement learning performance.
The expectation appearing in the policy gradient expression must be estimated in practice. To reduce gradient estimation variance, we use value function baselines in our experiments, and sample a meta-batch of $\bar{B}$ batched trajectories from $\mathbb{E}_{\bar{P}^{\phi^{i}}}$.
Relationship to prior shaping methods.
We now contrast our policy gradient algorithm to two closely related methods, M-FOS (Lu et al., 2022) and Shaper (Khan et al., 2024). Like COALA-PG, M-FOS is a model-free meta reinforcement learning method. Unlike the approached followed here, though, it aims to solve the bilevel co-player shaping problem of Eq. 1, treating meta- and inner-policy networks separately. Moreover, the M-FOS parameter update is not derived as the policy gradient on the batched co-player shaping POMDP introduced above, and current-episode returns are overemphasized compared to future-episode returns (see Appendix E). This leads to a biased parameter update, which results in learning inefficiencies. We comment on other existing bilevel shaping methods in Appendix D.

### v2

@@ Abstract
Self-interested individuals often fail to cooperate, posing a fundamental challenge for multi-agent learning. How can we achieve cooperation among self-interested, independent learning agents? Promising recent work has shown that in certain tasks cooperation can be established between “learning-aware" agents who model the learning dynamics of each other. Here, we present the first unbiased, higher-derivative-free policy gradient algorithm for learning-aware reinforcement learning, which takes into account that other agents are themselves learning through trial and error based on multiple noisy trials. We then leverage efficient sequence models to condition behavior on long observation histories that contain traces of the learning dynamics of other agents. Training long-context policies with our algorithm leads to cooperative behavior and high returns on standard social dilemmas, including a challenging environment where temporally-extended action coordination is required. Finally, we derive from the iterated prisoner’s dilemma a novel explanation for how and when cooperation arises among self-interested learning-aware agents.
@@ 1 Introduction
From self-driving autonomous vehicles to personalized assistants, there is a rising interest in developing agents that can learn to interact with humans (Collins et al., 2024; Gweon et al., 2023), and with each other (Park et al., 2023; Vezhnevets et al., 2023). However, multi-agent learning comes with significant challenges that are not present in more conventional single-agent paradigms. This is perhaps best seen through the study of “social dilemmas", general-sum games which model the tension between cooperation and competition in abstract form (von Neumann and Morgenstern, 1947). Without further assumptions, letting agents independently optimize their individual objectives on such games results in poor outcomes and a lack of cooperation (Tan, 1993; Claus and Boutilier, 1998).
@@
First, for general-sum games, reaching an equilibrium point does not necessarily imply appropriate behavior because there can be many sub-optimal equilibria (Fudenberg and Levine, 1998; Shoham and Leyton-Brown, 2008). Second, the control problem an agent faces is non-stationary from its own viewpoint, because other agents themselves simultaneously learn and adapt (Hernandez-Leal et al., 2017). Centralized training algorithms sidestep non-stationarity issues by sharing agent information (Sunehag et al., 2017), but this transformation into a global learning problem is usually prohibitively costly, and impossible to implement when agents must be developed separately (Zhang et al., 2021).
@@
The above two fundamental issues have hindered progress in multi-agent reinforcement learning, and have limited our understanding of how self-interested agents may reach high returns when faced with social dilemmas. In this paper, we join a promising line of work on “learning awareness" that has been shown to improve cooperation (Foerster et al., 2018a). The key idea behind such approaches is to take into account the learning dynamics of other agents explicitly, rendering it into a meta-learning problem (Schmidhuber, 1987; Bengio et al., 1990; Hochreiter et al., 2001).
@@
The present paper contains two main novel results on learning awareness in general-sum games. First, we introduce a new learning-aware reinforcement learning rule derived as a policy gradient estimator. Unlike existing methods (Foerster et al., 2018a, b; Xie et al., 2021; Balaguer et al., 2022; Lu et al., 2022; Willi et al., 2022; Cooijmans et al., 2023; Khan et al., 2024; Aghajohari et al., 2024), it has a number of desirable properties: (i) it does not require computing higher-order derivatives, (ii) it is provably unbiased, (iii) it can model minibatched learning algorithms, (iv) it is applicable to scalable architectures based on recurrent sequence policy models, and (v) it does not assume access to privileged information, such as the opponents’ policies or learning rules. Our policy gradient rule significantly outperforms previous model-free methods in the general-sum game setting. In particular, we show that efficient learning-aware learning suffices to reach cooperation in a challenging sequential social dilemma involving temporally extended actions (Leibo et al., 2017) that we adapt from the Melting Pot suite (Agapiou et al., 2023). Second, we analyze the iterated prisoner’s dilemma (IPD), a canonical model for studying cooperation among self-interested agents (Rapoport, 1974; Axelrod and Hamilton, 1981). Our analysis uncovers a novel mechanism for the emergence of cooperation through learning awareness, and explains why the seminal learning with opponent-learning awareness algorithm due to Foerster et al. (2018a) led to cooperation in the IPD.
@@ 2 Background and problem setup
We consider partially observable stochastic games (POSGs; Kuhn, 1953) consisting of a tuple
$\left(\mathcal{I},\mathcal{S},\mathcal{A},P_{t},P_{r},P_{i},\mathcal{O},P_{o},\gamma,T\right)$ with =$\mathcal{I}=\{1,\ldots,n\}$ a finite set of $n$ agents, $\mathcal{S}$ the state space, $\mathcal{A}=\times_{i\in\mathcal{I}}\mathcal{A}^{i}$ the joint action space, $P_{t}(S_{t+1}\mid S_{t},A_{t})$ the state transition distribution, $P_{i}(S_{0})$ the initial state distribution, $P_{r}=\times_{i\in\mathcal{I}}P_{r}^{i}(R\mid S,A)$ the joint factorized reward distribution with $R=\{R^{i}\}_{i\in\mathcal{I}}$ and bounded rewards $R^{i}$, $\mathcal{O}=\times_{i\in\mathcal{I}}\mathcal{O}^{i}$ the joint observation space, $P_{o}(O_{t}\mid S_{t},A_{t-1})$ the observation distribution, $\gamma$ the discount factor, $t$ the time step, and $T$ the horizon. We use superscript $i$ to indicate agent-specific actions, observations and rewards, $-i$ to indicate all agent indices except $i$, and we omit the superscript for joint actions, observations and rewards. As agents only receive partial state information, they benefit from conditioning their policies $\pi^{i}(a^{i}_{t}\mid x^{i}_{t};\phi^{i})$ on the observation history $x^{i}_{t}=\{o^{i}_{k}\}_{k=1}^{t}$ (Åström, 1965; Kaelbling et al., 1998), with $\phi^{i}$ the policy parameters. Note that the observations can contain the agent’s actions on previous timesteps.
@@ 2.1 General-sum games and their challenges
We focus on general-sum games, where each agent has their own reward function, possibly different from those of other agents. Specifically, we consider mixed-motive general-sum games that are neither zero-sum nor fully-cooperative. Analyzing and solving such general-sum games while letting every agent individually and independently maximize their rewards (a setting often referred to as “fully-decentralized reinforcement learning”; Albrecht et al., 2024) is a longstanding problem in the fields of machine learning and game theory for two primary reasons, described below.
Non-stationarity of the environment. In a general-sum game, each agent aims to maximize its expected return $J^{i}(\phi^{i},\phi^{-i})=\mathbb{E}_{P^{\phi^{i},\phi^{-i}}}\left[\sum_{t=1}^{T}\gamma^{t}R^{i}_{t}\right]$, with $P^{\phi^{i},\phi^{-i}}$ the distribution over environment trajectories $x_{T}$ induced by the environment dynamics, the policy $\pi^{i}(a^{i}\mid x^{i};\phi^{i})$ of agent $i$, and the policies $\pi^{-i}$ of all other agents. Importantly, the expected return $J^{i}(\phi^{i},\phi^{-i})$ does not only depend on the agent’s own policy, but also on the current policies of the other agents. As other agents are updating their policies through learning, the environment which includes the other agents is effectively non-stationary from a single agent’s perspective. Furthermore, the actions of an agent can influence this non-stationarity by changing the observation histories of other agents, on which they base their learning updates.
@@ 2.2 Co-player learning awareness
We aim to address the above two major challenges of multi-agent learning in this paper. Our work builds upon recent efforts that are based on adding a meta level to the multi-agent POSG, where the higher-order variable represents the learning algorithm used by each agent (Lu et al., 2022; Balaguer et al., 2022; Khan et al., 2024). In this meta-problem, the environment includes the learning dynamics of other agents. At the meta-level, one episode now extends across multiple episodes of actual game play, allowing the “ego agent", $i$, to observe how its co-players, $-i$, learn, see Fig. 1. The goal of this meta-agent may be intuitively understood as that of shaping co-player learning to its own advantage.
@@
Following Lu et al. (2022), we first introduce a meta-game with a single meta-agent whose goal is to shape the learning of naive co-players to its advantage. This meta-game is defined formally as a single-agent partially observable Markov decision process (POMDP) $(\tilde{\mathcal{S}},\tilde{\mathcal{A}},\tilde{P}_{t},\tilde{P}_{r},\tilde{P}_{i},\tilde{\mathcal{O}},\tilde{\gamma},M)$. The meta-state consists of the policy parameters $\phi^{-i}$ of all co-players together with the agent’s own parameters $\phi^{i}$. The meta-environment dynamics represent the fixed learning rules of the co-players, and the meta-reward distribution represents the expected return $J^{i}(\phi^{i},\phi^{-i})$ collected by agent $i$ during an inner episode, with “inner" referring to the actual game being played. The initialization distribution $\tilde{P}_{i}$ reflects the policy initializations of all players. Finally, we introduce a meta-policy $\mu(\phi^{i}_{m+1}\mid\phi^{i}_{m},\phi^{-i}_{m};\theta)$ parameterized by $\theta$, that decides the update to the parameter $\phi^{i}_{m+1}$ (the meta-action) to shape the co-player learning towards highly rewarding regions for agent $i$ over a horizon of $M$ meta steps. This leads to the co-player shaping problem
@@ 2.3 Single-level co-player shaping by leveraging sequence models
To proceed with this approach, we must first reformulate the meta-game. In particular, we must deal with a difficulty that is not present in single-agent meta reinforcement learning (e.g., Wang et al., 2016; Duan et al., 2017), which stems from the fact that co-players generally update their own policies based on multiple inner episodes (“minibatches"), without which reinforcement learning cannot practically make progress. Here, we solve this by defining the environment dynamics over $B$ parallel trajectories, with $B$ the size of the minibatch of inner episode histories that co-players use to update their policies at each inner episode boundary (see Fig. 1A).
Batched co-player shaping POMDP. We define the batched co-player shaping POMDP
@@
$(\bar{\mathcal{S}},\bar{\mathcal{A}},\bar{P}_{t},\bar{P}_{r},\bar{P}_{i},\bar{\mathcal{O}},\bar{\gamma},M,B)$, with hidden states consisting of the hidden environment states of the $B$ ongoing inner episodes, combined with the current parameters $\phi^{-i}_{m}$ of all co-players; environment dynamics $\bar{P}_{t}$ simulating $B$ environments in parallel, combined with updating the co-player’s policy parameters $\phi^{-i}$, and resetting the environments at each inner episode boundary; initial state distribution $\bar{P}_{i}$ that initializes the co-player policies and initializes the environments for the first inner episode batch; and finally, an ego-agent policy $\bar{\pi}^{i}(\bar{a}^{i}_{l}\mid\bar{h}^{i}_{l};\phi^{i})$ parameterized by $\phi^{i}$, which determines a distribution over the batched action $\bar{a}^{i}_{l}=\{a^{i,b}_{l}\}_{b=1}^{B}$, based on the batched long history $\bar{h}^{i}_{l}=\{h^{i,b}_{l}\}_{b=1}^{B}$. We refer to each element of the latter as a long history $h^{i,b}_{l}$, with long time index $l$ running across multiple episodes, from $l=1$ until $l=MT$. It should be contrasted to the inner episode history $x_{t}^{i}$, which runs from $t=1$ to $t=T$ and thus only reflects the current (inner) game history.
@@
The POMDP introduced above suggests using a sequence policy $\bar{\pi}^{i}(\bar{a}^{i}_{l}\mid\bar{h}^{i}_{l};\phi^{i})$ that is aware of the full minibatch of long histories and which produces a joint distribution over all current actions in the minibatch. However, as we aim to use our agents not only to shape naive learners, but also to play against/with each other, we require a policy that can be used both in a batch setting with naive learners, and in a single-trajectory setting with other learning-aware agents. Within our single-level approach, we achieve this by factorizing the batch-aware policy $\bar{\pi}^{i}(\bar{a}^{i}_{t}\mid\bar{h}^{i}_{l};\phi^{i})$ into $B$ independent policies with shared parameters $\phi^{i}$, $\bar{\pi}^{i}(\bar{a}^{i}_{l}\mid\bar{h}^{i}_{l};\phi^{i})=\prod_{b=1}^{B}\pi^{i}(a^{i,b}_{l}\mid h^{i,b}_{l};\phi^{i})$. Thanks to the batched POMDP, we can now pose co-player shaping as a standard (single-level, single-agent) expected return maximization problem:
@@ 3 Co-agent learning-aware policy gradients
3.1 A policy gradient for shaping naive learners
@@
$\nabla_{\phi^{i}}\bar{J}(\phi^{i})=\mathbb{E}_{\bar{P}^{\phi^{i}}}\left[\sum_{b=1}^{B}\sum_{l=1}^{MT}\nabla_{\phi}\log\pi^{i}(a^{i,b}_{l}\mid h^{i,b}_{l})\left(\frac{1}{B}\sum_{l^{\prime}=l}^{m_{l}T}r_{l^{\prime}}^{i,b}+\frac{1}{B}\sum_{b^{\prime}=1}^{B}\sum_{l^{\prime}=m_{l}T+1}^{MT}r_{l^{\prime}}^{i,b^{\prime}}\right)\right].$ (3)
We provide a proof in Appendix D. There are three important differences between COALA-PG and naively applying policy gradient methods to individual trajectories in a batch. (i) Each gradient term for an individual action $a_{l}^{i,b}$ takes into account the future inner episode returns averaged over the whole minibatch, instead of the future return along trajectory $b$ (see Fig. 2). This allows taking into account the influence of this action on the parameter update of the naive learner, which influences all trajectories in the minibatch after that update. (ii) Instead of averaging the policy gradients for each trajectory in the batch, COALA-PG accumulates (sums) them. This is important, as otherwise the learning signal would vanish in the limit of large minibatches. Intuitively, when a naive learner uses a large minibatch for its updates, the effect of a single action on the naive learner’s update is small ($O(\frac{1}{B})$), and this must be compensated by summing all such small effects. (iii) To ensure a correct balance between the return from the current inner episode $m_{l}$ and the return from future inner episodes, COALA-PG rescales the current episode return by $\frac{1}{B}$. Figure 12 in App. H shows empirically that COALA-PG correctly balances the policy gradient terms arising from the current inner episode return versus the future inner episode returns, whereas M-FOS (Lu et al., 2022) and a naive policy gradient that ignores the other parallel trajectories over-emphasize the current inner episode return, causing them to loose the co-player shaping learning signals. We will later show experimentally in Section 5 that correct treatment of minibatches critically affects reinforcement learning performance.
The expectation appearing in the policy gradient expression must be estimated. To reduce gradient estimation variance, we resort to standard practices, including generalized advantage estimation (Schulman et al., 2016) and sampling a meta-batch of $\bar{B}$ batched trajectories from $\mathbb{E}_{\bar{P}^{\phi^{i}}}$ (c.f. Appendix A).
Relationship to prior shaping methods. We now contrast our policy gradient algorithm to two closely related methods, M-FOS (Lu et al., 2022) and Shaper (Khan et al., 2024). Like COALA-PG, M-FOS is a model-free meta reinforcement learning method. Unlike the approached followed here, though, it aims to solve the bilevel co-player shaping problem of Eq. 1, treating meta- and inner-policy networks separately. Moreover, the M-FOS parameter update is not derived as the policy gradient on the batched co-player shaping POMDP introduced above, and current-episode returns are overemphasized compared to future-episode returns (see Appendix G). This leads to a biased parameter update, which results in learning inefficiencies. We comment on other existing bilevel shaping methods in Appendix F.